You are performing a right-tailed t-test with a sample size of 33

If α=.025 ,

find the critical value, to two decimal places.
Answer #1

Solution :

Given that,

sample size = n = 33

Degrees of freedom = df = n - 1 = 33 - 1 = 32

\alpha = 0.025

\alpha / 2 = 0.025 / 2 = 0.0125

t\alpha /2,df = t0.0125 , 32 = 2.348

Critical value = 2.348